# Politecnico di Torino, DIMEAS, Corso Duca degli Abruzzi, 24 10129 Torino, Italy

### v1

@@ IV.4.2 Long-term statistical prediction
Figure 19: Long-term statistical predictions for MFE model with classical and quantum reservoir computing.
Figure 19 presents the results for the statistical predictions against the true value. Both reservoirs can predict long-term statistics with high accuracy for the smaller kinetic energy values. Particularly, for $k<0.15$ the prediction agrees with the true value of the statistics. This analysis provides a statistical measure of the prediction of extreme events. Both quantum and classical reservoirs can predict most of the extreme events ($k<0.10$) in the test set. For the higher kinetic energy values, the log scale amplifies the difference between true value and predictions. The accuracy in the prediction of extreme events could be improved by considering a larger reservoir size, which scales better for QRC as compared to CRC, as shown previously in Figs. 17 and 18.
@@ V Conclusion
By exploiting quantum-computer ansatzë and entanglement, we analyse and design reservoir computers to predict chaotic dynamics and extreme events from data. First, we show that the design of a feature map, which encapsulates entanglement and data-encoding layers, is key to the reservoir’s performance. Second, we design the recurrence-free quantum reservoir computer (RF-QRC), which has a small circuit depth and does not have recurrent connections. This enables the training of the network without an additional classical feedback loop, which makes the implementation suitable to current noisy intermediate-scale quantum devices. Third, we numerically analyse quantum reservoir computers and the proposed RF-QRC on prototypical chaotic dynamical systems, from low- to higher-dimensional. We show that (a) the forecasting capability of state-of-the-art classical and quantum reservoir computers increases with the reservoir size until saturation; (b) for low-dimensional chaotic systems, classical reservoir computers have larger predictability than quantum reservoir computers for the same degrees of freedom; (c) for higher-dimensional chaotic systems, RF-QRC has larger predictability than classical reservoir computers for the same degrees of freedoms; and (d) RF-QRC can predict extreme events whilst scaling better than state-of-the-art reservoir computers. The RF-QRC requires smaller reservoir sizes as compared to classical reservoir computers (for the same performance), which significantly decreases the computational cost of matrix inversion in the training. The RF-QRC can be used to encode classical and quantum data on a quantum computer to make time-series predictions. This work opens new opportunities for the prediction of chaotic dynamics and extreme events with quantum reservoir computing. Current work is focused on the analysis of hardware and environmental noise, and finite-sampling errors.
@@ Acknowledgements.
Data Availability Statement
The data that support the findings of this study are available from the corresponding author upon reasonable request.
@@ Appendix A Quantum Circuits for different ansätze
The Figs. 20-23 represent the quantum circuits for different ansätze presented in Tab. 1. The rotation angles $X_{\Theta}$ represent the mapped classical data rescaled to the interval $[0,2\pi]$.
@@ Appendix B Memory of the Reservoir Networks
In this section, we discuss the empirical metrics including fading memory and linear Memory Capacity (MC) Boyd and Chua (1985); Carroll (2022); Farkaš et al. (2016); noa (b) for different reservoir architectures. In order to efficiently reconstruct the reservoir dynamics, reservoir computers must satisfy the Echo State Property (ESP) Jaeger (2001); Lukoševičius (2012). In classical reservoir networks, ESP is enforced by scaling the hyperparameter $\rho$ (Sec. II.1). In quantum reservoir networks, unitary evolution is norm-preserving, i.e. $(U^{{\dagger}}U=I)$, and therefore no tuneable hyperparameter $\rho$ exists. In both, quantum and classical reservoir computing, we perform a washout interval by discarding a limited number of initial reservoir states.
The Short-term Memory Capacity (MC) is a performance measure that is used to characterize the linear memory capacity of reservoir computers Jaeger (2002). MC quantifies the maximally possible linear correlations between the current reservoir states and previous input data, $\textbf{u}(t-d)$, and can be computed as
$MC=\sum_{d=1}^{d_{max}}MF_{d},\quad\quad{MF}_{d}=\frac{cov^{2}(\boldsymbol{u}_{d}(t),{\boldsymbol{u}(t-d)})}{\sigma^{2}(\boldsymbol{u}_{d}(t))\sigma^{2}({\boldsymbol{u}(t)})},$ (18)
@@
where $cov(\cdot,\cdot)$ is the covariance, $\sigma^{2}(\cdot)$ is the variance, $d$ is the delay, $\boldsymbol{u}(t-d)$ is the delayed prediction for a particular $d$ value and $\boldsymbol{u}_{d}(t)$ is the current true input value. Although the total memory capacity (MC) comprises infinitely many terms ($d_{max}=\infty$), for practical reasons it is sufficient to truncate the sum at a finite $d_{max}$. For a comparison of the memory capacities of CRC and QRC we choose $d_{max}=25$. While the MC quantifies the linear memory of the reservoir, the dynamical system itself has some correlations that can increase the MC. To quantify the memory of the reservoir itself, we train our quantum and classical reservoir networks on statistically independent (i.i.d) uniform distributions. In Figs. 24 and 25, we compare the linear memory capacity of the classical reservoir with QRC-C1 (architecture with recurrence) and QRC-C4 (recurrence-free architecture). The global MC is bounded by the size of the reservoir and is sensitive to the choice of hyperparameters Carroll (2022). We compute our MC for various hyperparameter values and reservoir sizes ($N=32,512$).
The results in Fig. 24 and 25 demonstrate that CRC exhibits the highest MC value followed by QRC-C1 and then QRC-C4. A high memory capacity in CRC highlights that the reservoir contains linear information about the input time series for a larger delay value $d$. In contrast, QRC-C4 does not have any recurrent or feedback loop. Due to the removal of the recurrence reservoir memory is lost. We emphasize that the higher MC value is not indicative of better predictability. This is because of two reasons: (a) The short-term memory capacity quantifies only the linear memory of the reservoir. (b) Maximizing the linear memory and nonlinear functional approximation are mutually exclusive and are referred to as the memory non-linearity trade-off Storm et al. (2022). We find comparable performances in short and long-term prediction tasks, irrespective of the MC value, for both classical and quantum reservoirs.

### v2

@@ IV.4.2 Long-term statistical prediction
Figure 20: MFE system. Long-term statistical predictions with classical and quantum reservoir computing.
Figure 20 presents the results for the statistical predictions against the true value. Both reservoirs can predict long-term statistics with high accuracy for the smaller kinetic energy values. Particularly, for $k<0.15$ the prediction agrees with the true value of the statistics. This analysis provides a statistical measure of the prediction of extreme events. Both quantum and classical reservoirs can predict most of the extreme events ($k<0.10$) in the test set. For the higher kinetic energy values, the log scale amplifies the difference between true value and predictions. The accuracy in the prediction of extreme events could be improved by considering a larger reservoir size, which scales better for QRC as compared to CRC, as shown previously in Figs. 18 and 19.
To further visualize the statistical predictions and compare CRC and RF-QRC, we predicted the velocity components of the flow fields from the predicted Fourier modes. In Fig. 21 the results of the average streamwise velocities are shown for the true data, and compared against classical and quantum reservoir predictions on the test data set. The predictions of RF-QRC approximate the true flow field better than the predictions from CRC, which have larger deviations. We conjecture that RF-QRC can be used to reconstruct flowfields with better accuracy than CRC with similar reservoir sizes. The prediction of flowfields with a better accuracy also allows the RF-QRC model to predict turbulent bursts of relaminarization and thus kinetic energies (extreme events) with better prediction horizon as previously shown in Figs. 18 and 19.
Figure 21: MFE system. The time average of the square of the midplane velocity along x. (a) True data (b) QRC-C4 predictions (c) CRC predictions for the best-performing networks with a reservoir size of 512 (9 qubits)
@@ V Conclusion
By exploiting quantum-computer ansätze and entanglement, we analyze and design reservoir computers to predict chaotic dynamics and extreme events from data. First, we show that the design of a feature map, which encapsulates entanglement and data-encoding layers, is key to the reservoir’s performance. Second, we design the Recurrence-Free Quantum Reservoir Computer (RF-QRC), which has a small circuit depth that does not scale with the number of qubits (reservoir size), unlike previous proposals of QRC. RF-QRC also does not have recurrent connections at each time step. This enables the training of the network without an additional classical feedback loop, which makes the implementation suitable to current noisy intermediate-scale quantum devices. Third, we numerically analyze quantum reservoir computers and the proposed RF-QRC on prototypical chaotic dynamical systems, from low- to higher-dimensional. We show that (a) the forecasting capability of state-of-the-art classical and quantum reservoir computers increases with the reservoir size until saturation; (b) for low-dimensional chaotic systems, classical reservoir computers have larger predictability than quantum reservoir computers for the same degrees of freedom; (c) for higher-dimensional chaotic systems, RF-QRC has larger predictability than classical reservoir computers for the same degrees of freedom; and (d) RF-QRC can predict extreme events while scaling better than state-of-the-art reservoir computers. The RF-QRC requires smaller reservoir sizes as compared to classical reservoir computers for the same performance. The RF-QRC can be used to encode classical and quantum data on a quantum computer to make time-series predictions. This work opens new opportunities for the prediction of chaotic dynamics and extreme events with quantum reservoir computing. Current work is focused on the analysis of hardware and environmental noise, and finite-sampling errors.
@@ Appendix A Quantum Circuits for different ansätze
The Figs. 22-25 represent the quantum circuits for different ansätze presented in Tab. 1. The rotation angles $X_{\Theta}$ represent the mapped classical data rescaled to the interval $[0,2\pi]$.
@@ Appendix B Memory of the Reservoir Networks
In this section, we discuss the empirical metrics including fading memory and linear Memory Capacity (MC) [58, 59, 60, 61, 62] for different reservoir architectures. In order to efficiently reconstruct the reservoir dynamics, reservoir computers must satisfy the Echo State Property (ESP) [15, 39]. In classical reservoir networks, ESP is enforced by scaling the hyperparameter $\rho$ (Sec. II.1). In quantum reservoir networks, unitary evolution is norm-preserving, i.e. $(U^{{\dagger}}U=I)$, and therefore no tuneable hyperparameter $\rho$ exists. In both, quantum and classical reservoir computing, we perform a washout interval by discarding a limited number of initial reservoir states.
The Short-term Memory Capacity (MC) is a performance measure that is used to characterize the linear memory capacity of reservoir computers [62]. MC quantifies the maximally possible linear correlations between the current reservoir states and previous input data, $\textbf{u}(t-d)$, and can be computed as
$MC=\sum_{d=1}^{d_{max}}MF_{d},\quad\quad{MF}_{d}=\frac{cov^{2}(\pmb{u}_{d}(t),{\pmb{u}(t-d)})}{\sigma^{2}(\pmb{u}_{d}(t))\sigma^{2}({\pmb{u}(t)})},$ (18)
@@
where $cov(\cdot,\cdot)$ is the covariance, $\sigma^{2}(\cdot)$ is the variance, $d$ is the delay, $\pmb{u}(t-d)$ is the delayed prediction for a particular $d$ value and $\pmb{u}_{d}(t)$ is the current true input value. Although the total memory capacity (MC) comprises infinitely many terms ($d_{max}=\infty$), for practical reasons it is sufficient to truncate the sum at a finite $d_{max}$. For a comparison of the memory capacities of CRC and QRC we choose $d_{max}=25$. While the MC quantifies the linear memory of the reservoir, the dynamical system itself has some correlations that can increase the MC. To quantify the memory of the reservoir itself, we train our quantum and classical reservoir networks on statistically independent (i.i.d) uniform distributions. In Figs. 26 and 27, we compare the linear memory capacity of the classical reservoir with QRC-C1 (architecture with recurrence) and QRC-C4 (recurrence-free architecture). The global MC is bounded by the size of the reservoir and is sensitive to the choice of hyperparameters [59]. We compute our MC for various hyperparameter values and reservoir sizes ($N=32,512$).
The results in Fig. 26 and 27 demonstrate that CRC exhibits the highest MC value followed by QRC-C1 and then QRC-C4. A high memory capacity in CRC highlights that the reservoir contains linear information about the input time series for a larger delay value $d$. In contrast, QRC-C4 does not have any recurrent or feedback loop. Due to the removal of the recurrence reservoir memory is lost. We emphasize that the higher MC value is not indicative of better predictability. This is because of two reasons: (a) The short-term memory capacity quantifies only the linear memory of the reservoir. (b) Maximizing the linear memory and nonlinear functional approximation are mutually exclusive and are referred to as the memory non-linearity trade-off [63]. We find comparable performances in short and long-term prediction tasks, irrespective of the MC value, for both classical and quantum reservoirs.